OpenStudy (anonymous):

factor the polynomial completely (y+4)^2 - 2(y+4) +1

OpenStudy (anonymous):

put y+4=x x^2-2x+1=x^2-x-x+1=x(x-1)-1(x-1)=(x-1)(x-1) replace the value of x